OpenStudy (anonymous):

this is confusing.... please help! Write the equation of the horizontal or vertical line: (3, -4), slope is undefined. y = 0 x = 3 x = -4 y = -4

OpenStudy (anonymous):

Which of those passes through x = 3 or y = -4?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

do you know how to graph?

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

then you know what a coordinate plane is, right?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i didnt learn that

OpenStudy (anonymous):

the coordinate plane is what the graph is called.

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

ok... so if you know how to graph do you know how to graph the line "x = -4"

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

there is no y intercept meaning it only passes through the x axis

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then which of those line equations pass through (3,-4)?

OpenStudy (anonymous):

x=3? i think...

OpenStudy (anonymous):

you should be down to 2 choices

OpenStudy (anonymous):

or y=0???

OpenStudy (anonymous):

y = 0 is incorrect

OpenStudy (anonymous):

so does that mean x=3 is correct?

OpenStudy (anonymous):

x = 3 is a possible choice, but I won't say if it's correct or not yet

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there is one more choice that you need to rule out

OpenStudy (anonymous):

before you can be certain

OpenStudy (anonymous):

well idk it can u just tell me it

OpenStudy (anonymous):

if x = 3 and that is a possible choice then obviously the only other choice would be the one that passes through y = -4

OpenStudy (anonymous):

now, do you know what undefined means?

OpenStudy (anonymous):

kinda ... not really

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

if a slope is 0 then it is horizontal

OpenStudy (anonymous):

if a slope is undefined it is...?

OpenStudy (anonymous):

vertical?

OpenStudy (anonymous):

ding ding ding... now which one of the equations is a vertical line? :)

OpenStudy (anonymous):

x = 3 or y = -4

OpenStudy (anonymous):

y=-4

OpenStudy (anonymous):

nope, y = -4 is horizontal

OpenStudy (anonymous):

oh....

OpenStudy (anonymous):

because it only passes through the y axis

OpenStudy (anonymous):

so its x=3

OpenStudy (anonymous):

Yurp. :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now you can solve all of those questions with ease!

OpenStudy (anonymous):

can u help me with my next questions???

OpenStudy (anonymous):

"questions"?

OpenStudy (anonymous):

its alot different

OpenStudy (anonymous):

i mean question

OpenStudy (anonymous):

I guess so.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Find the equation of a line parallel to the given line: y = -3/4x -1. y = -3/4x + 2 y = 3/4x – 1 y = 4/3x – 1 y = 3/4x + 1

OpenStudy (anonymous):

this one is really easy to solve... just look for the equation that has the same slope, but a different y intercept.

OpenStudy (anonymous):

there is only one equation in there that has the same slope. :)

OpenStudy (anonymous):

ohhh so its y=-3/4x+2

OpenStudy (anonymous):

ding ding ding

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

btw, slope intercept form is y = mx + b b = y intercept

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

ok this is my last question....

OpenStudy (anonymous):

Write an equation in slope-intercept form for an equation of a line perpendicular to y = 1/2x that intersects at point (-2, 3). y = -2x – 1 y = 2x + 7 y = 2x + 3 y = -2x +1